OpenStudy (anonymous):

Find the margin error when p=83% and n=300

OpenStudy (nurali):

In absence of specific information , 95% confidence level can be assumed 1) Margin of error = Confidence coefficient*Standard error of p Confidence coefficient is the critical value of z for 95% confidence level = 1.96 Standard error of p = sqrt [p*(1-p)/n] = sqrt [0.83*0.17/300] = 0.0217 Margin of error = 1.96*0.0217 = 0.04